OpenStudy (anonymous):

(3) Sketch the region and find the area of the region bounded by y = x 5 − x, y = 0, 0 6 x 6 2.

OpenStudy (leader):

Could you please clarify. Just restate it but please put enters between equations.

OpenStudy (anonymous):

sure, basically sketch a graph using these two functions y=x^5-x and y=0. Then after that find the area between both functions bounded by 0<x<2

OpenStudy (leader):

Ok so the equations are: y=x^(5)-x and y=0 and 0<x<2

OpenStudy (leader):

Do yu have to do it by yourself or is the use of a graphing calculator permitted?

OpenStudy (anonymous):

not graphing calculator permitted

OpenStudy (anonymous):

no*

OpenStudy (leader):

ok

OpenStudy (leader):

first let's graph y=x^5-x

OpenStudy (leader):

How many x-intercepts should it have?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

i'd say a minimum of 2

OpenStudy (leader):

Nevermind that was a bad question.

OpenStudy (leader):

So it is odd powered so the range is (-infinity,infinity)

OpenStudy (leader):